I need to get the solution to these equations using the substitution method:
-x+y=7 and -1/2-y=2

Answer by stanbon(75887)

get the solution to these equations using the substitution method:
-x+y=7 and -1/2-y=2
---------------------
y = -1/2 - 2
y = -5/2
-----
Substitute and solve for "x":
-x -(5/2) = 7
-x = 7+(5/2)
-x = 9.5
x = -9.5
========================
Cheers,
Stan H.
========================